# Draw-In-Mind: Rebalancing Designer-Painter Roles in Unified Multimodal Models Benefits Image Editing

### v2

@@ 5 Conclusion
In this paper, we identify a crucial issue in existing image editing models, i.e., imbalanced division of responsibilities, where the generator is burdened with complex reasoning, leading to reduced performance. To address this, we propose the Draw-In-Mind (DIM) dataset, consisting of two parts: (i) DIM-T2I, 14M web-crawled image–text pairs with carefully crafted long-context prompts that provide a foundation for complex CoT comprehension in editing; and (ii) DIM-Edit, 233K high-quality image editing pairs with detailed and precise CoT imagination. By training on the DIM dataset and incorporating an external designer during editing, we present DIM-4.6B-Edit, which achieves SOTA or competitive performance on ImgEdit and GEdit-Bench-EN while maintaining a tiny overall and trainable parameter size. These results validate our motivation to shift the design responsibility from the generation module to the understanding module, as well as the high efficiency of our proposed CoT-guided DIM dataset.
@@ Appendix A Additional Experiments
Detailed Performance on GEdit-Bench-EN. Table 8 and 9 summarize overall and detailed task-wise performance of different models on GEdit-Bench-EN, respectively. Our DIM-4.6B-Edit ranks just behind the in-domain tester, i.e., Step1X-Edit, while surpassing all other out-of-domain competitors. Moreover, among out-of-domain testers, DIM-4.6B-Edit is the only model that consistently preserves both semantic consistency and perceptual quality. This demonstrates the effectiveness of DIM-Edit, where edits with high perceptual fidelity are precisely aligned with CoT-style imagination, thereby ensuring semantic correctness.
@@
Visualizations of Different Editing Operations. Beyond Figure 3 in the manuscript, we further visualize the edits of Janus-4o, Step1X-Edit, and our DIM-4.6B-Edit under the operations of add, change, remove, replace, and style transfer in Figure 4567, and 8, respectively. As shown, DIM-4.6B-Edit consistently preserves the overall layout while performing natural edits. For instance, in Figure 4, Janus-4o fails to generate details of the wooden cabin, while Step1X-Edit places the chimney on the river, which is counterfactual. In contrast, our DIM-4.6B-Edit carefully adds the wooden cabin while ensuring naturalness. In Figure 5, Janus-4o fails to follow the color change instruction. Step1X-Edit changes the singer’s shirt to blue but also alters fine details such as the collar shape. By comparison, our DIM-4.6B-Edit changes the shirt to red while preserving all details, including the shadow cast by the hand. In Figure 6, both DIM-4.6B-Edit and Step1X-Edit perform successful removals, whereas Janus-4o fails to remove the seaplane. In Figure 7, only DIM-4.6B-Edit captures the semantics of “majestically” and generates a roaring lion. Finally, in Figure 8, although all three models succeed in style transfer, only DIM-4.6B-Edit captures subtle visual cues, such as the green grass in the last row, and repaints them faithfully in the edits.
Appendix B DIM-T2I Analysis Dimensions
Figure 9 and 10 illustrate the 21 analysis dimensions and their corresponding prompts used in the DIM-T2I annotation process. These dimensions were carefully selected to encompass both broad world knowledge and fine-grained visual attributes, such as illumination and shadow effects, which are crucial for high-quality image generation. As a result, DIM-T2I provides not only semantically rich supervision but also consistent aesthetic guidance, forming a strong foundation for precise image editing. This is further demonstrated by the trained DIM-4.6B-T2I model, which faithfully follows user instructions at the semantic level while producing perceptually high-quality images under limited parameter budgets, achieving high GenEval scores and low FID on MJHQ-30K.
@@ Appendix C DIM-Edit Data Collection Pipeline
  • UltraEdit (Zhao et al., 2024). To ensure editing consistency, we retain only those edit pairs that satisfy the following conditions: (i) the CLIP similarity between the source and edited images is greater than 0.9; (ii) the DINOv2 similarity is greater than 0.9; (iii) the SSIM score is greater than 0.8; and (iv) the prompt does not contain “rainbow”, since many edit pairs meeting (i)–(iii) are associated with low-quality “rainbow” edits. After filtering, we obtain roughly 160K edit pairs.
@@ Appendix D DIM-Edit Quality Assessment
Figure 11: The percentage distribution of each quality level in DIM-Edit.

### v3

@@ 5 Conclusion
In this paper, we identify a crucial issue in existing image editing models, i.e., imbalanced division of responsibilities, where the generator is burdened with complex reasoning, leading to reduced performance. To address this, we propose the Draw-In-Mind (DIM) dataset, consisting of two parts: (i) DIM-T2I, 14M web-crawled image-text pairs with carefully crafted long-context prompts that provide a foundation for complex CoT comprehension in editing; and (ii) DIM-Edit, 233K high-quality image editing pairs with detailed and precise CoT imagination. By training on the DIM dataset and incorporating an external designer during editing, we present DIM-4.6B-Edit, which achieves SOTA or competitive performance on ImgEdit and GEdit-Bench-EN while maintaining a tiny overall and trainable parameter size. These results validate our motivation to shift the design responsibility from the generation module to the understanding module, as well as the high efficiency of our proposed CoT-guided DIM dataset.
@@ Appendix A Additional Experiments
Detailed Performance on GEdit-Bench-EN. Table 11 and 12 summarize overall and detailed task-wise performance of different models on GEdit-Bench-EN, respectively. Our DIM-4.6B-Edit ranks just behind the in-domain tester, i.e., Step1X-Edit, while surpassing all other out-of-domain competitors. Moreover, among out-of-domain testers, DIM-4.6B-Edit is the only model that consistently preserves both semantic consistency and perceptual quality. This demonstrates the effectiveness of DIM-Edit, where edits with high perceptual fidelity are precisely aligned with CoT-style imagination, thereby ensuring semantic correctness.
@@
Table 14: The ImgEdit performance of different models with/without using DIM CoT as instruction.
Model Params CoT Add Adjust Extract Replace Remove Background Style Hybrid Action Overall
DIM-4.6B-Edit Und[Uncaptioned image] $|$ Gen[Uncaptioned image] 3.53 3.23 2.01 3.49 1.47 3.42 4.79 2.35 3.64 3.10
4.09 3.47 2.30 4.00 3.43 3.87 4.92 2.85 4.08 3.67
Janus-4o Und[Uncaptioned image] $|$ Gen[Uncaptioned image] 3.35 3.35 2.25 3.01 2.18 3.32 4.71 2.49 4.04 3.19
3.95 2.74 2.49 3.59 2.28 3.31 4.72 2.62 4.02 3.30
Step1X-Edit Und[Uncaptioned image] $|$ Gen[Uncaptioned image] 3.88 3.14 1.76 3.40 2.41 3.16 4.63 2.64 2.52 3.06
3.56 2.47 1.81 3.13 2.02 2.84 4.18 1.80 2.46 2.70
Impact of DIM CoT for Different Models. To investigate the impact of DIM-style CoT on different models, we evaluated the performance of Janus-4o and Step1X-Edit when directly provided with the same CoT blueprints as input instructions on ImgEdit. The results are presented in Table 14. Based on these results, we have the following observations and analysis:
  • DIM-4.6B-Edit is explicitly trained on complex CoT-style blueprints from the DIM-Edit dataset, it achieves superior CoT comprehension. Consequently, it demonstrates substantial performance gains when DIM-style CoTs are applied during inference.
  • Janus-4o employs an end-to-end fine-tuning approach, which minimizes the gap between instruction understanding and generation. This makes it more robust to input distribution shifts. While it possesses mild CoT comprehension capabilities and benefits slightly from DIM-style CoTs, the performance gain is less pronounced compared to DIM-4.6B-Edit.
  • Step1X-Edit adopts a training recipe similar to ours (using a frozen understanding core), this design makes it susceptible to input distribution shifts when facing unseen instruction formats. It struggles to process CoT inputs effectively, leading to performance degradation when DIM-style CoTs are applied.
Based on these findings, we conclude that superior CoT comprehension is pivotal for enhancing editing performance. This finding validates our strategy of fostering CoT comprehension by constructing DIM-T2I and utilizing DIM-4.6B-T2I as the initialization for the editing task.
Table 15: The performance of DIM-4.6B-T2I/Edit on understanding benchmarks.
Model Params MME-P MMB SEED MMMU $M$-Vet
Janus (Wu et al., 2025a) 1.3B[Uncaptioned image] 1338.0 69.4 63.7 30.5 34.3
Emu3-Gen (Wang et al., 2024b) 8.0B[Uncaptioned image] - 58.5 68.2 31.6 37.2
Show-o (Xie et al., 2024) 1.3B[Uncaptioned image] 1097.2 - - 26.7 -
Show-o2-7B (Xie et al., 2025b) 7.0B[Uncaptioned image] 1620.5 79.3 69.8 48.9 -
Janus-Pro-7B (Chen et al., 2025c) 7.0B[Uncaptioned image] 1567.1 79.2 72.1 41.0 50.0
BAGEL (Deng et al., 2025) 14.0B[Uncaptioned image] 1687.0 85.0 - 55.3 67.2
MetaQuery-L (Pan et al., 2025) 3.0B[Uncaptioned image] $|$ 3.2B[Uncaptioned image] 1574.3 78.6 73.8 53.1 63.2
DIM-4.6B-T2I/Edit 3.0B[Uncaptioned image] $|$ 1.6B[Uncaptioned image] 1574.3 78.6 73.8 53.1 63.2
Understanding Performance. Since the MLLM component is frozen during DIM training, its understanding performance remains unaffected and is identical to the results reported in the original paper. To ensure clarity regarding the model’s capabilities, we report the corresponding understanding performance in Table 15. Our experiments demonstrate that DIM-4.6B-Edit achieves superior editing performance even when utilizing a relatively small MLLM under a frozen setting. This finding highlights the flexibility of our approach: users can seamlessly upgrade to advanced MLLMs to unlock even greater understanding and editing performance. Such integration is straightforward, as our streamlined architecture and training recipe avoid the need for intricate parameter tuning.
Appendix B Additional Visualizations
B.1 Visualization of Different Editing Operations.
Beyond Figure 3 in the manuscript, we further visualize the edits of Janus-4o, Step1X-Edit, and our DIM-4.6B-Edit under the operations of add, change, remove, replace, and style transfer in Figure 4567, and 8, respectively. As shown, DIM-4.6B-Edit consistently preserves the overall layout while performing natural edits. For instance, in Figure 4, Janus-4o fails to generate details of the wooden cabin, while Step1X-Edit places the chimney on the river, which is counterfactual. In contrast, our DIM-4.6B-Edit carefully adds the wooden cabin while ensuring naturalness. In Figure 5, Janus-4o fails to follow the color change instruction. Step1X-Edit changes the singer’s shirt to blue but also alters fine details such as the collar shape. By comparison, our DIM-4.6B-Edit changes the shirt to red while preserving all details, including the shadow cast by the hand. In Figure 6, both DIM-4.6B-Edit and Step1X-Edit perform successful removals, whereas Janus-4o fails to remove the seaplane. In Figure 7, only DIM-4.6B-Edit captures the semantics of “majestically” and generates a roaring lion. Finally, in Figure 8, although all three models succeed in style transfer, only DIM-4.6B-Edit captures subtle visual cues, such as the green grass in the last row, and repaints them faithfully in the edits.
@@ Appendix B Additional Visualizations
B.2 Visualization of the Draw-In-Mind Workflow’s Impact on Image Editing.
Relying solely on numerical metrics may not intuitively convey the practical impact of the Draw-In-Mind workflow on image generation. To address this, we present Figure 9101112, and 13 to showcase several advanced usage scenarios. These examples demonstrate complex cases that are successfully handled by DIM-Edit-4.6B, highlighting capabilities that remain beyond the reach of current baseline methods.
Instruction Disambiguation. In Figure 9, the user instruction presents an inherent ambiguity due to the presence of three lemons on the table. This task necessitates precise multi-object localization and removal, which is a challenge that proves difficult without the Draw-In-Mind paradigm, as standard models often struggle with the required multi-object reasoning. Consequently, both the 7B Janus-4o and 12B Step1X-Edit fail to execute the edit correctly. Similarly, when CoT is disabled, our DIM-4.6B-Edit also fails to remove all targets. However, with DIM CoT enabled, the generated design blueprints effectively disambiguate the instruction. They accurately localize the three lemons to the right of the vase and ensure their complete removal, while perfectly preserving the integrity of the unedited regions.
Edit Navigation and Structural Planning. In Figure 10, the user instruction presents two distinct challenges: (i) determining the optimal placement for a wooden cabin, and (ii) identifying the appropriate structural integration for a chimney. These dual requirements impose a significant burden on the generation model. Consequently, in Janus-4o’s output, the chimney is nearly invisible, while Step1X-Edit places the cabin counterintuitively close to the river. Similarly, DIM without CoT fails to simultaneously resolve the cabin placement and chimney addition. In contrast, DIM powered by CoT effectively navigates these challenges. It observes that ”the trees thin out on the right side” (GLP) and selects this area as the optimal location (EAL). It then explicitly envisions the cabin’s appearance, including a chimney emitting smoke (EII), ultimately yielding the most plausible and high-quality edit among all competitors.
Commonsense-guarded Editing. In Figure 11, the editing task presents a subtle complexity: it requires commonsense reasoning regarding scale. From the same viewpoint, a cat should appear significantly smaller than a horse. All baseline models, including our own DIM w/o CoT, overlook this physical constraint, simply replacing the horse with a cat of identical dimensions. In contrast, DIM with CoT successfully leverages commonsense reasoning. It recognizes the size discrepancy and executes a ”commonsense-guarded” edit, placing a naturally scaled cat at the target location, thereby preserving scene realism.
Advanced Causal Editing. In Figure  12, we present an advanced causal editing scenario where the instruction implies the target quantity (referencing ”the second prime number”) rather than stating it explicitly. Unsurprisingly, all baseline models fail to resolve this implicit requirement. In contrast, DIM with CoT swiftly infers the correct number of cherries and executes a successful edit, demonstrating its ability to handle knowledge-intensive instructions.
Advanced Temporal Editing. Figure 13 illustrates the most complex temporal editing scenario, which necessitates a deep understanding of chemical reaction dynamics. Similar to the previous example, none of the baseline models succeed in this task. In contrast, DIM with CoT accurately characterizes the reaction process and executes physically plausible edits, demonstrating its capability to handle sophisticated temporal reasoning.
Figure 9: The edits of Janus-4o, Step1X-Edit, DIM w/o CoT, and DIM-4.6B-Edit when the user instruction is ambiguous. DIM CoT is capable of instruction disambiguation under this case.
Figure 10: The edits of Janus-4o, Step1X-Edit, DIM w/o CoT, and DIM-4.6B-Edit when the user instruction requires localization and involves fine-grained structure modification. DIM CoT is capable of edit navigation and structural planning under this case.
Figure 11: The edits of Janus-4o, Step1X-Edit, DIM w/o CoT, and DIM-4.6B-Edit when the user instruction involves implicit commonsense constraint. DIM CoT is capable of commonsense-guarded editing under this case.
Figure 12: The edits of Janus-4o, Step1X-Edit, DIM w/o CoT, and DIM-4.6B-Edit when the user instruction requires causal reasoning. DIM CoT is capable of advanced causal editing under this case.
Figure 13: The edits of Janus-4o, Step1X-Edit, DIM w/o CoT, and DIM-4.6B-Edit when the user instruction requires temporal reasoning. DIM CoT is capable of advanced temporal editing under this case.
B.3 Visualization of Failure Cases
We are also open to discuss the limitations of our work, and provide three failure types with six specific cases in Figure 1415, and 16 to intuitively show the boundaries of our DIM-4.6B-Edit.
Large-scale All-in-One Editing. In Figure 14, the instructions involve simultaneous multi-step edits, a task that remains essentially challenging for almost all editing models, and one where DIM-Edit also encounters difficulties.
  • For the first case, Janus-4o and Step1X-Edit completely fail to follow the physical laws dictating that the wooden tower should collapse. Our DIM-4.6B-Edit successfully imitates a scene of imminent collapse; however, it fails to preserve the exact appearance of the individual wooden blocks, as too many objects are involved in the manipulation.
  • For the second case, Janus-4o and Step1X-Edit fail to change the view at all. While our DIM-4.6B-Edit completes the primary editing task, some fine-grained details are distorted (e.g., the window of the shoreside house is missing).
Text and Logic Editing. In Figure 15, where instructions involve complex text rendering and logical editing, DIM-4.6B-Edit struggles due to a combination of data scarcity and inherent VAE compression issues.
  • For the first case, the use of SANA1.5’s VAE with a 32x downsampling rate makes complex text rendering particularly challenging, a difficulty exacerbated by the lack of targeted training data. In contrast, Step1X-Edit employs an 8x downsampling VAE and is trained on proprietary, text-specific in-house data, allowing it to perform relatively well. We regard this as a necessary trade-off between efficiency and rendering quality: as shown in Table 10, DIM-4.6B-Edit requires only 6 seconds to complete an edit with a 200+ word CoT, whereas Step1X-Edit takes 28 seconds with a short raw prompt.
  • For the second case, all editing models fail. This is fundamentally because none of the models, including DIM-4.6B-Edit, are specifically trained on geometric data. The underlying painter struggles to even draw these shapes, let alone edit them. We believe crafting such datasets remains a valuable and under-explored topic for future research.
Reference-free Editing (in Pixel Space). In Figure 16, the reference image does not provide a strong pixel constraint for the target image. Consequently, this task resembles multimodal generation rather than strict editing. All models fail here because existing editing architectures typically enforce strong pixel alignment with the source image.
  • For the first case, which requests a view of the Golden Gate Bridge, Janus-4o and Step1X-Edit are completely ineffective. DIM-4.6B-Edit struggles to break free from the structural constraints of the reference image, resulting in a ”scratchy” and distorted view that fails to meet the objective.
  • For the second case, where the task involves a re-imagination of the source scene, Janus-4o produces a black-and-white edit, and Step1X-Edit fails completely. DIM-4.6B-Edit generates the most plausible result, successfully covering the scene with white snow. However, because the transformation fundamentally alters the source structure, specific details such as the castle are inevitably distorted.
In summary, the majority of failure cases arise when the task necessitates either generating an image that diverges drastically from the source or rendering complex text and geometric shapes. Even in these challenging scenarios, DIM-4.6B-Edit demonstrates superior instruction-following capabilities compared to baseline models. These limitations highlight persistent challenges within the current landscape of open-source editing models. We suggest that future research directions, such as intelligent routing that dynamically selects between T2I generation and editing pipelines based on instruction intent, offer promising avenues for resolving these issues, though significant progress is still required in the field.
Figure 14: The edits of Janus-4o, Step1X-Edit, and DIM-4.6B-Edit for the failure type large-scale all-in-one editing.
Figure 15: The edits of Janus-4o, Step1X-Edit, and DIM-4.6B-Edit for the failure type text and logic editing.
Figure 16: The edits of Janus-4o, Step1X-Edit, and DIM-4.6B-Edit for the failure type reference-free editing (in pixel space).
@@ Appendix C DIM-Edit Data Collection Pipeline
  • UltraEdit (Zhao et al., 2024). In addition to the prompt quality evaluation and optimization in the DIM-Edit creation pipeline (Figure 2), which aligns textual prompts with actual editing behaviors, we employ three image-to-image metrics on the UltraEdit dataset to improve visual consistency and stabilize training: (i) CLIP image similarity, (ii) DINOv2 similarity, and (iii) SSIM. These metrics are used to select edit pairs that maintain consistent visual appearances. We retain only those edit pairs that satisfy the following conditions: (i) the CLIP similarity between the source and edited images is greater than 0.9; (ii) the DINOv2 similarity is greater than 0.9; (iii) the SSIM score is greater than 0.8; and (iv) the prompt does not contain “rainbow”, since many edit pairs meeting (i)–(iii) are associated with low-quality “rainbow” edits. After filtering, we obtain roughly 160K edit pairs.
@@ Appendix D DIM-Edit Quality Assessment
Figure 17: The percentage distribution of each quality level in DIM-Edit judged by GPT-4.1.
@@
Figure 18: (a) The win rate of the optimized DIM-Edit instruction and the raw instruction. (b) The quality of the Edited Image Imagination (EII).
We further conducted a human verification study. Specifically, we randomly sampled 25 instances from each of the data sources listed in Appendix C, resulting in a comprehensive evaluation set of 100 samples. Three human annotators were then recruited to assess the quality of the CoTs from two distinct perspectives:
Evaluation of Optimized Instructions (Start of CoT). We presented annotators with both the raw instructions and the optimized instructions from DIM-Edit, alongside the corresponding source-edit image pairs. Annotators were tasked with selecting the instruction that best reflected the actual editing operations. A ”Tied” option was included for cases where neither instruction was sufficiently accurate. The metric reported is the average win rate for each instruction type.
Evaluation of Edited Image Imagination (End of CoT). We asked annotators to assess the alignment between the Edited Image Imagination (EII) and the actual edited image. The quality was categorized into three levels: (i) Perfect Match, (ii) Minor Errors, and (iii) Major Errors. The metric reported is the percentage distribution across these error levels.
This efficient evaluation protocol enables a rapid yet robust assessment of the overall CoT quality within DIM-Edit. The results for both the instruction optimization (Win Rate) and the Edited Image Imagination (Error Distribution) are summarized in Figure 18, in which we have the following analysis:
  • Consistency with MLLM Assessment. These results align closely with the MLLM-based quality assessment presented in Appendix D, where over 80% of DIM-Edit CoTs were judged clearer than the raw instructions, with no factual errors detected. Even in ”Tied” cases where the optimization was not deemed strictly superior, the semantics of the raw instruction were fully preserved, ensuring that the optimization process introduces no regression.
  • Analysis of Minor Errors. We observed that minor errors typically relate to subtle environmental inconsistencies, such as slight shifts in brightness (e.g., ”the image should be a bit lighter”). These artifacts usually stem from the VAE’s inability to perfectly reconstruct raw images in AI-generated pairs (e.g., from UltraEdit), leading to a slight loss of high-frequency features. As these discrepancies are barely perceptible to the human eye, they have a negligible impact on overall training efficiency.
  • Analysis of Major Errors. Instances classified as having major errors generally correspond to extremely challenging scenarios where the edits are minute (e.g., the removed object occupies less than 2% of the pixels). These cases are difficult even for human annotators and advanced MLLMs like GPT-4o. Given their extreme rarity, these outliers do not adversely affect the stability of the training procedure.
Overall, the CoTs produced by our DIM-Edit pipeline maintain high quality and serve as effective design blueprints. This high data quality directly translates to better editing capabilities, as evidenced by the superior performance of the DIM-4.6B-Edit model trained on this dataset.
Appendix E DIM-T2I Analysis Dimensions
Figure 19 and 20 illustrate the 21 analysis dimensions and their corresponding prompts used in the DIM-T2I annotation process. The 21 dimensions were derived from a thorough literature review and an empirical analysis of existing understanding datasets and benchmarks. They are listed as follows:
MME (Fu et al., 2025), MMMU (Yue et al., 2024), MMMU-Pro (Yue et al., 2025), MMLU (Hendrycks et al., 2020), MMStar (Chen et al., 2024b), MMT-Bench (Ying et al., 2024), MM-Vet (Yu et al., 2023), MM-Vet V2 (Yu et al., 2024), LLaVA-Bench-Wild (Liu et al., 2023a), LLaVA-Bench-Wilder (Li et al., 2024a), WildVision (Lu et al., 2024), COCO (Lin et al., 2014), VQAv2 (Goyal et al., 2017), OK-VQA (Marino et al., 2019), TextCaps (Sidorov et al., 2020), TextVQA (Singh et al., 2019), AI2D (Kembhavi et al., 2016), ChartQA (Masry et al., 2022), DocVQA (Mathew et al., 2021), MathVista (Lu et al., 2023), MIA-Bench (Qian et al., 2024), MegaBench (Chen et al., 2024a), RWQA (xAI, 2024), OCRBench (Liu et al., 2023b), GSM8K (Cobbe et al., 2021), GPQA (Rein et al., 2024), IFEval (Zhou et al., 2023).
We believe that the aspects emphasized in widely recognized understanding datasets and benchmarks effectively capture the most frequent interactions between humans and objects in the real world. This makes them an ideal foundation for learning text-to-image generation tasks involving long and complex instructions. By constructing prompts that span these diverse fields, DIM-4.6B-T2I not only masters long-form instruction processing but also acquires the broad world knowledge necessary to facilitate sophisticated CoT comprehension and precise editing, thereby achieving high GenEval scores and low FID on MJHQ-30K.